# Explanation User Interfaces: A Systematic Literature Review

### v1

@@ Abstract.
Artificial Intelligence (AI) is one of the major technological advancements of this century, bearing incredible potential for users through AI-powered applications and tools in numerous domains. Being often black-box (i.e., its decision-making process is unintelligible), developers typically resort to eXplainable Artificial Intelligence (XAI) techniques to interpret the behaviour of AI models to produce systems that are transparent, fair, reliable, and trustworthy. However, presenting explanations to the user is not trivial and is often left as a secondary aspect of the system’s design process, leading to AI systems that are not useful to end-users. This paper presents a Systematic Literature Review on Explanation User Interfaces (XUIs) to gain a deeper understanding of the solutions and design guidelines employed in the academic literature to effectively present explanations to users. To improve the contribution and real-world impact of this survey, we also present a framework for Human-cEnteRed developMent of Explainable user interfaceS (HERMES) to guide practitioners and academics in the design and evaluation of XUIs.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: SUBMITTEDpublicationmonth: 5price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Human-centered computing User interface designccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recent years have witnessed a remarkable interest in the field of Artificial Intelligence (AI), which has become a prominent topic in both academia and industry. The rapid advancements in AI technologies have led to their widespread adoption across various sectors, including healthcare, finance, and transportation. The machine learning subfield of AI has gained particular attention due to its ability to provide discriminative and predictive capabilities, enabling the development of sophisticated applications such as image recognition, natural language processing, and autonomous systems. The increasing complexity of machine learning models has raised concerns about their interpretability and transparency, in particular for those models that are based on deep learning architectures. These models often operate as “black boxes,” making it challenging for users to understand their decision-making processes (Guidotti et al., 2019). This has led to a growing demand for eXplainable Artificial Intelligence (XAI) techniques that can elucidate the decision-making processes of these models.
Black-box machine learning models pose various threats to their users. For example, the risk of unfair and unjust bias increases when the model’s inner workings are not clear (Mehrabi et al., 2022). Therefore, worldwide legislative bodies are attempting to regulate the use of such systems. For example, the European Union (EU) recently released the AI act with this goal (European Parliament and Council of the European Union, 2024). Most of these regulations advocate the need for AI systems to be accountable, fair, and explainable (European Parliament and Council of the European Union, 2024). In this context, XAI may be beneficial.
@@
XAI is a sub-field of AI, initially formalised by the Defense Advanced Research Projects Agency (DARPA) in 2017 (Gunning and Aha, 2019), which focuses on “opening” black-box models, making their inner workings clear, explainable, and interpretable (Guidotti et al., 2019; Gunning and Aha, 2019). Although different techniques for explaining black-box models exist (as will be detailed in Section 2.1), a common and precise definition of “good” explanations has yet to be established. Several studies have pointed out that the majority of the work on XAI is based on researchers’ intuition of what qualifies as a “good” explanation (Miller, 2019; Abdul et al., 2018; Liao and Varshney, 2021; Lipton, 2018; Cheng et al., 2019), effectively framing XAI primarily as an algorithmic problem. As a result, there is a gap between XAI algorithms designed by researchers and their deployment in real-word scenarios  (Liao and Varshney, 2021). To fill this gap, Human-Computer Interaction (HCI) research has recently begun to focus on AI-enabled applications and XAI, as testified by the birth of the research field of Human-Centered Artificial Intelligence (HCAI), which reframes AI-development in a human-centred perspective by focusing on users’ needs (Shneiderman, 2022).
Since the first conceptualisation of XAI by DARPA, the design of XAI systems followed two different stages: the design of the algorithmic model itself, and the design of the User Interface (UI) (Chromik and Butz, 2021). Such UIs are usually referred to as Explanation User Interfaces (Chromik and Butz, 2021). As with other types of UIs, XUIs must be designed following a human-centred approach to be successful and usable (ISO, 2018).
@@
In this study, we present a Systematic Literature Review (SLR), following Kitchenham’s procedure (Kitchenham, 2004), that aims at providing a comprehensive overview of the design of XUIs. Our overall goal is to identify design trends among human-centred XUIs and XAI algorithms to provide a cohesive understanding of the Human-Centered eXplainable Artificial Intelligence (HCXAI) landscape. More precisely, we contribute to the state-of-the-art systematising the knowledge on four main interrelated aspects of XUIs design; (i) The XAI models used to generate the explanations, (ii) The solutions adopted to provide the explanations, (iii) The user studies and human-centred techniques used to evaluate the designs, (iv) The design guidelines and frameworks that can help designers and practitioners.
This article is structured as follows: Section 2 presents key background concepts; Section 3 presents the methodology used to conduct the SLR and the four research questions that guided it; Sections 4, 5, 6 and 7 report the results to answer each of the four research questions; Section 8 presents a set of conclusions drawn by analyzing the dimensions of each RQ; moreover, it presents HERMES, a framework to guide practitioners in the design and evaluation of XUIs; Section 9 reports open research challenges; Section 10 presents the limitations of our study, highlighting any potential threat to its validity; finally, Section 11 concludes the article.
@@ 2. Background
This section introduces the background of our work. It provides an overview of the field of XAI, reprising a classification of methods already available in the literature that can guide in understanding the results of this SLR. Furthermore, it explores the need for a human-centred process and its implications for XUIs.
@@ 2.1. Explainable Artificial Intelligence
The increasing use of complex and opaque AI models, such as deep learning architectures, has resulted in the emergence of “black-box” systems. The decision-making processes of such systems are often difficult to interpret, even for AI engineers. XAI confronts this issue by offering techniques that enhance the understandability of AI model predictions for humans (Guidotti et al., 2019). This increased clarity fosters trust and accountability (Shin, 2021; Ferrario and Loi, 2022), while also mitigating the risks associated with automated decision-making in high-stakes situations.
In XAI, explanations can be classified as either global or local. Global explanations offer insights into how a black-box model operates, aiding users in understanding its overall functioning (Guidotti et al., 2019). In contrast, local explanations focus on understanding the model’s predictions for individual cases (Guidotti et al., 2019). Additionally, the type of data processed by the AI model also influences explanations. XAI methods are most commonly applied to tabular, text, and image data, but they can also handle other types of data such as time series (Theissler et al., 2022), audio (Akman and Schuller, 2024), and graphs (Nandan et al., 2025). A possible classification of XAI techniques, based on the type of explanation used, is the following (Guidotti et al., 2019; Adadi and Berrada, 2018):
@@
A small set of interpretable models (e.g., decision trees, decision rules, and linear models) can be used to explain more complex models by approximating their global or local behaviour. For example, a decision tree can approximate a complex model behaviour, allowing its interpretation as a chain of if-then-else rules.
@@
Explanations may be provided by describing each feature’s weight (i.e., importance) in the decision-making process. Such techniques are referred to as “posthoc” since they can be employed after the complex model has been built. Examples of such techniques are Locally Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016) and SHapley Additive exPlanation (SHAP) (Lundberg and Lee, 2017). The main difference between the two approaches is that while LIME can only provide local explanations, SHAP can also produce global explanations. Both techniques are model-agnostic since they can be applied to any existing model.
@@
Saliency masks highlight subsets of the original instance (e.g., set of pixels of an image or words in a sentence) mainly responsible for a certain outcome. Thus, a “salient mask” visually highlights the aspects of the input instance determining the outcome. Saliency masks are generally used to explain deep neural networks; e.g., Class Activation Mapping (CAM) (Zhou et al., 2016), and its generalisation Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020), highlight the contents of an image based on an attention mechanism, which is visualised as an explanation. Saliency masks are also helpful for text: rationales (Lei et al., 2016) are a short and coherent piece of text (e.g., a sentence) that represents a sufficient subset of words that can be used to predict the original text on their own.
@@
Prototype-based Explanations:
@@ 2.2. Towards Human-Centered Explainable Artificial Intelligence
Although XAI aims at addressing the challenges posed by AI-based systems, which, despite their high precision, frequently remain confined to research settings, it often falls short of producing usable systems that present their outcomes and explanations in a reliable, safe, and trustworthy manner. In real-world contexts, humans are generally unwilling or unable to trust black-box systems that do not provide insight into their decision-making processes (Combi et al., 2022). This lack of transparency prevents end-users from questioning AI decisions, potentially allowing harmful biases to go unnoticed (Stahl et al., 2023).
In response to these challenges, a new field of study has emerged in recent years at the intersection of HCI and AI, namely HCAI (Desolda et al., 2024b). HCAI proposes a new perspective on the interaction with AI, aiming to augment rather than replace humans and their expertise (Shneiderman, 2020). HCAI systems are designed to be ethically aligned, reflect human intelligence, and consider human factors (Xu, 2019). Recent research in HCAI suggests the need for a truly human-centered process in the design of AI systems, showing how there is no “one-size-fits-all” approach when designing AI systems, as different user goals bring the need for a different level of automation and control in a system (Esposito et al., 2024). One of the milestones of HCAI is the adoption of HCI methods for the design and development of HCAI. In particular, HCAI stresses the importance of user studies in eliciting requirements and validating final systems.
Human-Centered Design (HCD) is the general model adopted in HCI for the design of systems that satisfy users’ needs and expectations; it specifies that users are involved from the very beginning of the planning stage, and identifying user requirements becomes a crucial phase (Rogers et al., 2023). Various HCD-based design processes are now available, such as UCD Sprint (Larusdottir et al., 2022). HCD requires understanding who will use the system, where, and how to do what. The system is then designed by iterating a design-evaluation cycle. Being the design based on empirical knowledge of user behaviour, needs, and expectations, it is possible to avoid serious mistakes and to save re-implementation time to correct such mistakes. In HCAI, adopting HCD is necessary when a direct interaction between users and the system is expected (e.g., in consumer applications, educational software, or health-related support).
In this context, well-designed human-centred Human-AI interfaces are essential for the success of an AI system (Shneiderman, 2022; Raees et al., 2024; Desolda et al., 2024a; Esposito et al., 2024). Furthermore, when an AI system is deployed in high-risk domains like medicine, explainability is crucial (Gerdes, 2024; the Precise4Q consortium et al., 2020; Combi et al., 2022). Therefore, a thoughtfully designed, human-centered XUI is necessary.
@@ 2.3. Explanation User Interfaces
The notion of an interface mediating between the XAI algorithm and the user was initially introduced by DARPA (Gunning and Aha, 2019). DARPA divides the XAI process into two different stages: the generation of the explanations themselves, and the presentation of such explanations through UIs. The definition of this two-stage process is shared in the academic literature. Danilevsky et al. differentiate between explanation techniques and explanation visualizations: the former involves the generation of ‘rough’ explanations (usually propounded by AI researchers), while the latter concerns the ways ‘rough’ explanations can be presented to users (Danilevsky et al., 2020).
A rigorous definition of XUIs is formulated by Chromik and Butz, who define a XUI as ¡¡the sum of outputs of an XAI system that the user can directly interact with. An XUI may tap into the machine learning (ML) model or may use one or more explanation-generating algorithms to provide relevant insights for a particular audience¿¿ (Chromik and Butz, 2021). In general, XUIs can be designed for various forms of explanation, both local and global, and can be presented in textual or visual form, or a combination of the two  (Bove et al., 2023). XUIs can also leverage other modalities such as sound and tangible interfaces (Schneider et al., 2021b; Colley et al., 2022; Schuller et al., 2021).
According to the definition of XUIs (Chromik and Butz, 2021), two key aspects are fundamental to their development: interactivity and adaptability for different types of users. Interactivity, in particular, plays a central role in the explanation process (Miller, 2019). Research in HCI highlights the importance of designing interfaces that allow users to explore explanations freely (Abdul et al., 2018), ask follow-up questions, and access additional details as needed (Krause et al., 2016). Achieving this level of usability in real-world scenarios requires continuous user involvement throughout the engineering process, from the initial system design to the final testing phases (Rogers et al., 2023). However, despite its emphasis in HCI, user-centred design remains relatively underexplored in XAI and XUIs, highlighting a critical gap in the field. XUIs fall into the category of interfaces designed for AI-infused systems. For these interfaces, traditional UI principles remain essential; however, the unique and dynamic behaviour of AI systems necessitates the development of new guidelines tailored to their specific characteristics (Amershi et al., 2019).

### v2

@@ Abstract.
Artificial Intelligence (AI) is one of the major technological advancements of this century, bearing incredible potential for users through AI-powered applications and tools in numerous domains. Being often black-box (i.e., its decision-making process is unintelligible), developers typically resort to eXplainable Artificial Intelligence (XAI) techniques to interpret the behaviour of AI models to produce systems that are transparent, fair, reliable, and trustworthy. However, presenting explanations to the user is not trivial and is often left as a secondary aspect of the system’s design process, leading to AI systems that are not useful to end-users. This paper presents a Systematic Literature Review on Explanation User Interfaces (XUIs) to gain a deeper understanding of the solutions and design guidelines employed in the academic literature to effectively present explanations to users. To improve the contribution and real-world impact of this survey, we also present a platform to support Human-cEnteRed developMent of Explainable user interfaceS (HERMES) and guide practitioners and scholars in the design and evaluation of XUIs.
@@
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Human-centered computing User interface designccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recent years have witnessed a remarkable interest in the field of Artificial Intelligence (AI), which has become a prominent topic in both academia and industry. The rapid advancements in AI technologies have led to their widespread adoption across various sectors, including healthcare, finance, and transportation. The Machine Learning subfield of AI has gained particular attention due to its ability to provide discriminative and predictive capabilities, enabling the development of sophisticated applications such as image recognition, natural language processing, and autonomous systems. The increasing complexity of Machine Learning models has raised concerns about their interpretability and transparency, particularly for those models based on deep learning architectures. These models often operate as “black boxes,” making it challenging for users to understand their decision-making processes (Guidotti et al., 2019). This has led to a growing demand for eXplainable Artificial Intelligence (XAI) techniques that can elucidate the decision-making processes of these models. Black-box Machine Learning models pose various threats to their users. For example, the risk of unfair and unjust bias increases when the model’s inner workings are not clear (Mehrabi et al., 2022). Therefore, worldwide legislative bodies are attempting to regulate the use of such systems. For example, the European Union (EU) recently released the AI act with this goal (European Parliament and Council of the European Union, 2024). Most of these regulations advocate the need for AI systems to be accountable, fair, and explainable (European Parliament and Council of the European Union, 2024). In this context, XAI may be beneficial. XAI is a sub-field of AI, initially formalised by the Defense Advanced Research Projects Agency (DARPA) in 2017 (Gunning and Aha, 2019), which focuses on “opening” black-box models, making their inner workings clear, explainable, and interpretable (Guidotti et al., 2019; Gunning and Aha, 2019). Although different techniques for explaining black-box models exist (as will be detailed in Section 2.1), a common and precise definition of “good” explanations has yet to be established. Several studies have pointed out that the majority of the work on XAI is based on researchers’ intuition of what qualifies as a “good” explanation (Miller, 2019; Abdul et al., 2018; Liao and Varshney, 2021; Lipton, 2018; Cheng et al., 2019), effectively framing XAI primarily as an algorithmic problem. As a result, there is a gap between XAI algorithms designed by researchers and their deployment in real-word scenarios  (Liao and Varshney, 2021). To fill this gap, Human-Computer Interaction (HCI) research has recently begun to focus on AI-enabled applications and XAI, as testified by the birth of the research field of Human-Centered Artificial Intelligence (HCAI), which reframes AI-development in a human-centred perspective by focusing on users’ needs (Shneiderman, 2022). Since the first conceptualisation of XAI by DARPA, the design of XAI systems followed two different stages: the design of the algorithmic model itself, and the design of the User Interface (UI) (Chromik and Butz, 2021). Such UIs are usually referred to as Explanation User Interfaces (Chromik and Butz, 2021). As with other types of UIs, XUIs must be designed following a human-centred approach to be successful and usable (ISO, 2018a).
@@
In this study, we present a Systematic Literature Review (SLR), following Kitchenham’s procedure (Kitchenham, 2004), that aims at providing a comprehensive overview of the design of XUIs. Our overall goal is to identify design trends among human-centred XUIs and XAI algorithms to provide a cohesive understanding of the Human-Centered eXplainable Artificial Intelligence (HCXAI) landscape. More precisely, we contribute to the state-of-the-art systematising the knowledge on four main interrelated aspects of XUIs design; (i) The XAI models used to generate the explanations, (ii) The solutions adopted to provide the explanations, (iii) The user studies and human-centred techniques used to evaluate the designs, (iv) The design guidelines and frameworks that can help designers and practitioners. This article is structured as follows: Section 2 presents key background concepts; Section 3 presents the methodology used to conduct the SLR and the four research questions that guided it; Sections 4.1, 4.2, 4.3 and 4.4report the results to answer each of the four research questions; Section 5 presents a set of conclusions drawn by analyzing the dimensions of each RQ; moreover, it presents HERMES, a platform to guide practitioners in the design and evaluation of XUIs; Section 6 reports open research challenges; Section 7 presents the limitations of our study, highlighting any potential threat to its validity; finally, Section 8 concludes the article.
@@ 2. Background
This section provides an overview of the background of our work. It provides an overview of the field of XAI, reprising a classification of methods already available in the literature that can guide in understanding the results of this SLR. Furthermore, it explores the need for a human-centred process and its implications for XUIs.
@@ 2.1. Explainable Artificial Intelligence
[…]
@@
A small set of interpretable models, such as decision trees and decision rules, can be used to explain more complex models by approximating their global or local behaviour. For example, a decision tree can approximate the behaviour of a complex model, allowing its interpretation as a chain of if-then-else rules. This line of research builds upon foundational work in extracting symbolic rules from trained artificial neural networks (Andrews et al., 1995) and neuro-fuzzy systems (Mitra and Hayashi, 2000), recently systematized in reviews on symbolic knowledge extraction and injection (Ciatto et al., 2024).
@@
Explanations may be provided by describing each feature’s weight (i.e., importance) in the decision-making process. Such techniques are referred to as “posthoc” since they can be employed after the complex model has been built. An example of such techniques is Locally Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016), a technique to explain individual predictions by approximating the local behavior of a model with a simpler, interpretable model around a specific instance.
Shapley Values:
A particular instance of explanations based on feature importance is SHapley Additive exPlanation (SHAP), a technique based on cooperative game theory which uses Shapley values to attribute a model’s output to each input feature, ensuring consistency and local accuracy in explanations (Lundberg and Lee, 2017). Compared to an approach like LIME, which can only provide local explanations, SHAP can also produce global explanations, at the cost of a higher computational expense.
@@
Saliency masks highlight subsets of the original instance (e.g., set of pixels of an image or words in a sentence) mainly responsible for a certain outcome. Thus, a “salient mask” visually highlights the aspects of the input instance determining the outcome. Saliency masks are generally used to explain deep neural networks. E.g., Class Activation Mapping (CAM) (Zhou et al., 2016), and its generalisation Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020), highlight the contents of an image based on an attention mechanism, which is visualised as an explanation. Saliency masks are also helpful for text: rationales (Lei et al., 2016) are a short and coherent piece of text (e.g., a sentence) that represents a sufficient subset of words that can be used to predict the original text on their own.
@@
Exemplar-based/Prototype-based Explanations:
@@
Other:
Other approaches include emerging or ad-hoc techniques that do not fall among the traditional categories. Examples of these are Explicit Factor Model for explainable recommendations (Zhang et al., 2014), Visual Interactive Model Explorer for ML models debugging in sequential decision-making (Antar et al., 2024), approaches that exploit attention mechanisms for explaining deep neural networks (El Houda Dehimi and Tolba, 2024), and algorithmic-based (Desolda et al., 2023) or agent-based explanations (Weitz et al., 2021).
@@ 2.2. Towards Human-Centered Explainable Artificial Intelligence
Although XAI aims to address the challenges posed by AI-based systems, which, despite their high precision, often remain confined to research settings, it frequently falls short of producing usable systems that present their outcomes and explanations in a reliable, safe, and trustworthy manner. In real-world contexts, humans are generally unwilling or unable to trust black-box systems that do not provide insight into their decision-making processes (Combi et al., 2022). This lack of transparency prevents end-users from questioning AI decisions, potentially allowing harmful biases to go unnoticed (Stahl et al., 2023). In response to these challenges, a new field of study has emerged in recent years at the intersection of HCI and AI, namely HCAI (Desolda et al., 2024b). HCAI proposes a new perspective on the interaction with AI, aiming to augment rather than replace humans and their expertise (Shneiderman, 2020). HCAI systems are designed to be ethically aligned, reflect human intelligence, and consider human factors (Xu, 2019). Recent research in HCAI suggests the need for a truly human-centered process in the design of AI systems, showing how there is no “one-size-fits-all” approach when designing AI systems, as different user goals bring the need for a different level of automation and control in a system (Esposito et al., 2024). One of the milestones of HCAI is the adoption of HCI methods for the design and development of HCAI. In particular, HCAI stresses the importance of user studies in eliciting requirements and validating final systems. Human-Centered Design (HCD) is the general model adopted in HCI for the design of systems that satisfy users’ needs and expectations; it specifies that users are involved from the very beginning of the planning stage, and identifying user requirements becomes a crucial phase (Rogers et al., 2023). Various HCD-based design processes are now available, such as UCD Sprint (Larusdottir et al., 2022). HCD requires understanding who will use the system, where, and how to do what. The system is then designed by iterating a design-evaluation cycle. Being design-based on empirical knowledge of user behavior, needs, and expectations, it is possible to avoid serious mistakes and save reimplementation time to correct such mistakes. In HCAI, adopting HCD is necessary when a direct interaction between users and the system is expected (e.g., in consumer applications, educational software, or health-related support). In this context, well-designed human-centred Human-AI interfaces are essential for the success of an AI system (Shneiderman, 2022; Raees et al., 2024; Desolda et al., 2024a; Esposito et al., 2024). Furthermore, when an AI system is deployed in high-risk domains like medicine, explainability is crucial (Gerdes, 2024; the Precise4Q consortium et al., 2020; Combi et al., 2022). Therefore, a thoughtfully designed, human-centered XUI is necessary.
@@ 2.3. Explanation User Interfaces
The notion of an interface mediating between the XAI algorithm and the user was initially introduced by DARPA (Gunning and Aha, 2019). DARPA divides the XAI process into two different stages: the generation of the explanations themselves, and the presentation of such explanations through UIs. The definition of this two-stage process is shared in the academic literature. Danilevsky et al. differentiate between explanation techniques and explanation visualizations: the former involves the generation of ‘rough’ explanations (usually propounded by AI researchers), while the latter concerns the ways ‘rough’ explanations can be presented to users (Danilevsky et al., 2020). A rigorous definition of XUIs is formulated by Chromik and Butz, who define a XUI as ¡¡the sum of outputs of an XAI system that the user can directly interact with. An XUI may tap into the ML model or may use one or more explanation-generating algorithms to provide relevant insights for a particular audience¿¿ (Chromik and Butz, 2021). In general, XUIs can be designed for various forms of explanation, both local and global, and can be presented in textual or visual form, or a combination of the two  (Bove et al., 2023). XUIs can also leverage other modalities such as sound and tangible interfaces (Schneider et al., 2021a; Colley et al., 2022; Schuller et al., 2021).
According to the definition of XUIs (Chromik and Butz, 2021), two key aspects are fundamental to their development: interactivity and adaptability for different types of users. Adaptive systems change their behavior automatically, driven by context-aware mechanisms, including models of their users and of specific tasks. Adaptive systems are important because there is no ’typical’ user; there are many different users, and an individual user’s requirements can change over time (Fischer, 2023).
Interactivity plays a central role in the explanation process (Miller, 2019). Research in HCI highlights the importance of designing interfaces that allow users to explore explanations freely (Abdul et al., 2018), ask follow-up questions, and access additional details as needed (Krause et al., 2016). Achieving this level of usability in real-world scenarios requires continuous user involvement throughout the engineering process, from the initial system design to the final testing phases (Rogers et al., 2023). However, despite its emphasis in HCI, user-centred design remains relatively underexplored in XAI and XUIs, highlighting a critical gap in the field. XUIs fall into the category of interfaces designed for AI-infused systems. For these interfaces, traditional UI principles remain essential; however, the unique and dynamic behaviour of AI systems necessitates the development of new guidelines tailored to their specific characteristics (Amershi et al., 2019).